A series RLC circuit with R = 6 ohm, C = 0.02 Farad and L = 0.1 has no applied voltage. Find the subsequent current in the circuit if the initial charge, on the capacitor is q0 and the initial current is zero.
1
Expert's answer
2020-03-24T10:27:10-0400

Kirchhoff's voltage law:

"u_R+u_L+u_C=0"

where "u_R,u_L,u_C" are the voltage across "R,L,C" respectively.

Substituting in the constitutive equations:

"Ri(t)+L\frac{di(t)}{dt}+\frac{1}{C}\int \limits_{-\infty}^{t}i(\tau)d\tau=0"

Differentiating and dividing by "L" :

"\frac{d^2i(t)}{dt^2}+\frac{R}{L}\frac{di(t)}{dt}+\frac{1}{LC}i(t)=0"

This can usefully be expressed in a more generally applicable form: 

"\frac{d^2i(t)}{dt^2}+2 \alpha\frac{di(t)}{dt}+\omega^2_0i(t)=0\\
\alpha=\frac{R}{2L}, \omega_0=\frac{1}{\sqrt{LC}}"

The differential equation has the characteristic equation:

"s^2+2\alpha s+\omega_0^2=0"

The roots of the equation in "s" are: 

"s_{1,2}=-\alpha \pm\sqrt{\alpha^2-\omega_0^2}"

The general solution of the differential equation is an

exponential in either root or a linear superposition of both

"i(t)=Ae^{s_1t}+Be^{s_2t}"

The initial current is zero:

"i(0)=Ae^{0}+Be^{0}=0"

Therefore: "A=-B"  

"i(t)=A(e^{s_1t}-e^{s_2t})=\\
=Ae^{-\alpha t}(e^{\sqrt{\alpha^2-\omega_0^2}t}-e^{-\sqrt{\alpha^2-\omega_0^2}t})\\
\sqrt{\alpha^2-\omega_0^2}=\sqrt{(30)^2-10\cdot 50}=\\
=\sqrt{400}=20, \alpha=30"

therefore

"i(t)=A(e^{-10t}-e^{-50t})"

The initial charge on the capacitor is "q_0" and initial current is zero:

"L\frac{di(t)}{dt}|_{t=0}+\frac{q_0}{C}=0\\
\frac{di(t)}{dt}|_{t=0}=-500q_0\\
\frac{di(t)}{dt}=A(-10e^{-10t}+50e^{-50t)}\\
\frac{di(t)}{dt}|_{t=0}=40A\\
40A=-500q_0\\
A=-12.5q_0"


Therefore

"i(t)=-12.5q_0(e^{-10t}-e^{-50t})"